A. Transform 4x-8y=13 to slope intercept form
B. What is the slope of the line perpendicular to the one in A if it goes through the point (-11,17)?

Answer by TimothyLamb(4379)

L1:
4x - 8y = 13
---
8y = 4x - 13
y = (4/8)x - (13/8)
---
L1 in slope-intercept form:
y = (1/2)x - (13/8)
---
What is the slope of the line L2 perpendicular to line L1?
---
perpendicular lines have negative reciprocal slopes, so:
slope of L2 = -1/(1/2)
---
slope of L2 = -2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php